Circular No. 6045

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

JUPITER AND COMET 1993e P. Galopeau, E. Gérard, and A. Lecacheux, Observatoire de Paris,

Meudon, communicate: “The jovian synchrotron emission has been monitored with the Nançay Radio telescope at 1404, 1664, and 3263 MHz during three periods: Apr. 1–17, May 1–17, and July 9–28. The observations are made daily near 18^{h} UT and will last until Aug. 23. A large flux-density increase at all frequencies started on July 18, reached a maximum between July 20 and 22, and has decreased since then; the flux-density maximum was 30–40 percent higher (depending on frequency) than during the April and May runs, which showed no variation larger than 5 percent. Thus, there is no doubt that this outburst is due to the encounter of comet 1993e with Jupiter.”

SUPERNOVA 1968D IN NGC 6946 S. D. Van Dyk, University of California, Berkeley; S. D. Hyman, Sweet

Briar College; R. A. Sramek, National Radio Astronomy Observatory; and, K. W. Weiler, Naval Research Laboratory, report the radio recovery of the type-II supernova 1968D in the nearby spiral galaxy NGC 6946: “A candidate source was first found in the summed Very Large Array (VLA) data collected over several years for the monitoring of the radio SN 1980K in this galaxy at 20 and 6 cm (1.43 and 4.86 GHz, respectively; see Weiler et al. 1986, Ap.J. 301, 790; Weiler et al. 1992, Ap.J. 398, 248). Detection of the source has been confirmed by additional observations with the VLA in configuration A, made in 1994 April at these same wavelengths. SN 1968D has radio position α = 20^{h}33^{m}55^{s}.298, δ = +59^{o}59′09′′.07 (equinox 1950.0). The position of the radio nucleus of NGC 6946 has end figures 49^{s}.245, 58′49′′.24, giving a nuclear offset for SN 1968D of 45′′.3 east and 19′′.8 north, which is in excellent agreement with the optical offset position obtained near the time of discovery (IAUC 2057; Barbon et al. 1989, A.Ap. Suppl. 81, 421). The supernova has flux densities 0.15 ± 0.03 and 0.52 ± 0.03 mJy at 6 and 20 cm, respectively. The source is therefore nonthermal, with a steep spectral index α = −1.02 ± 0.15.”

COMET NAKAMURA-NISHIMURA-MACHHOLZ (1994m) Total visual magnitude estimates by C. E. Spratt, Victoria, BC (0.20-m

reflector): July 15.29 UT, 8.9; 28.28, 8.8; Aug. 1.27, 8.3.

1994 August 2 Daniel W. E. Green